the perimeter of a garden is 34 feet. I fthe lenght is 3 less than 3 times the width, what is the area in square feet of the garden
Answer by macston(5194)

the perimeter of a garden is 34 feet. If the length is 3 less than 3 times the width, what is the area in square feet of the garden
Width=W: Length=L=3W-3; Perimeter=P=2(L+W)
P=2(L+W)
34 ft=2(L+W) Divide each side by 2.
17 ft=L+W Substitute for L.
17 ft=(3W-3)+W Add 3 to each side.
20 Ft=4W Divide each side by 4.
5 ft=W The width of the garden is 5 ft.
L=3W-3=3(5 ft)-3=15 ft-3=12 ft The garden is 12 ft long.
Area=L*W=12 ft*5 ft=60 sq ft ANSWER: The area of the garden is 60 sq ft.